Circular No. 6343

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) R. M. West, European Southern Observatory (ESO), writes: “A series

of forty-four 2-sec exposures in B, V, and R was obtained under excellent conditions (seeing 0′′.7–0′′.8) on Mar. 19.33–19.40 UT with the ESO 3.5-m New Technology Telescope (+ SUSI camera; observers O. Hainaut and R. West). A comparison of two series of twelve R-frames each, at mean epochs Mar. 19.342 and 19.363 (i.e., a 30-min interval), by means of averaged, radially normalized frames, shows substantial morphological changes (on a scale of arcseconds) of the dust jets within 10′′ from the peak of light. In addition to a well-indicated clockwise rotation of ∼ 20^{o}, there are also changes apparent in the jet structure. There is also an unchanged straight ion tail at p.a. 276^{o}, noticeable at 0′′.7 (115 km projected) from the peak of light in both frames. These images have been placed on the ESO World Wide Web home-

page (http://www.eso.org/educnpubrelns/comet-hyakutake.html).”

Naked-eye m_{1}, coma-diameter, and tail-length estimates: Mar. 17.35 UT, 3.4, —, > 6^{o} (J. V. Scotti, Kitt Peak, AZ); 18.58, 2.5, 50′, 10^{o} (P. Camilleri, Cobram, Victoria, Australia); 19.22, 2.6, 1^{o}.1, — (J. E. Bortle, Stormville, NY); 19.34, 2.0, ∼ 80′, 20^{o} (C. S. Morris, Whitaker Peak, CA); 19.38, 2.5, 1^{o}.2, 15^{o} (A. Hale, Cloudcroft, NM); 19.73, 2.5, —, — (A. Beresford, Adelaide, South Australia).

GRO J1744–28 L. Bouchet, J. Paul, M. Gilfanov, and R. Sunyaev, on behalf of the

SIGMA/Granat team (Centre d’Etude Spatiale des Rayonnements, Toulouse; Service d’Astrophysique du CEA, Saclay; Space Research Institute, Moscow), report: “The SIGMA telescope onboard the Granat satellite started a new campaign of observing the galactic-center region on Mar. 15.62 UT, continuing until Mar. 29. In the first observation (effective duration 30 000 s), GRO J1744–28 was detected in the images up to ∼ 75 keV at α = 17^{h}41^{m}22^{s}, δ = −28^{o}45′.6 (equinox 1950.0; uncertainty ± 3′), only 1′.5 from the reported XTE position (IAUC 6309). Preliminary analysis gave the total (variable + continuous) flux of ∼ 150 mCrab in the band 40–75 keV, with a very soft spectrum. Also, 1E 1740.7–2942 was detected (well separated from GRO J1744–28) at the ∼ 3σ level in the same band. The two sources GRO J1744–28 and 1E 1740.7–2942 were also detected in a subsequent observation (duration 80 000 s) beginning on Mar. 17, which has shown a decline of the flux of GRO J1744–28 by a factor of ∼ 1.5 in the energy range 40–75 keV.”

1996 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green